# Action Reasoning Models that can Reason in Space

### v1

@@ 5.6 Steerability of MolmoAct
Evaluation Setups and Baselines. We aim to evaluate MolmoAct’s ability to steer robot actions, particularly when initial language instructions are ambiguous. Specifically, we investigate the effectiveness of different interaction mediums in guiding MolmoAct toward user-intended targets during task execution. For this purpose, we set up a pick_up_the_bowl task, training MolmoAct and the baseline model ($\pi_{0}$-FAST) with 100 collected demonstrations, each annotated with two distinct language instructions: one specifying the clean bowl and the other the dirty bowl, as depicted in Figure 7. During evaluation, we first provide ambiguous instructions such as "pick up the bowl," prompting MolmoAct to predict an initial trajectory towards one of the bowls. Subsequently, we test two steering methods: visual trace sketches to visually instruct the model toward the alternative bowl, and open-ended natural language instructions provided interactively by participants (N=10). For comparison, we also attempt to steer the actions of $\pi_{0}$-FAST by changing language instructions at test-time. Each model is evaluated in 15 trials, and the performance is evaluated according to the progression of the task.
Evaluation Results. Based on our experiments, we observed that MolmoAct is notably more steerable via visual trace inputs, achieving a success rate of 75%. Additionally, steering using visual traces significantly outperforms steering via open-ended natural language instructions by a margin of 33%. Lastly, we demonstrate that MolmoAct exhibits superior instruction-following capabilities compared to the baseline model, $\pi_{0}$-FAST. Specifically, when steering robot actions using language instructions, MolmoAct surpasses $\pi_{0}$-FAST by a substantial margin of 29%, highlighting its enhanced responsiveness to user commands.
@@ 6.1 Generalist robot manipulation policies
Recent advances in robotic manipulation have shifted from training narrow, single-task specialists to learning from large, diverse datasets spanning many scenes, tasks, and embodiments(Berscheid et al., 2019; Brohan et al., 2022; Dasari et al., 2019; Ebert et al., 2021; Fang et al., 2023; Jang et al., 2022; Khazatsky et al., 2024; Mandlekar et al., 2018; Walke et al., 2023; Shafiullah et al., 2023). This shift has enabled policies that not only excel within their training distribution but also generalize to out-of-distribution scenes, environments, language instructions, and novel objects (Pumacay et al., 2024; Xie et al., 2024; Lin et al., 2024). Much of this progress has been fueled by Large Language Models (LLMs) (O’Neill et al., 2024; Achiam et al., 2023; Groeneveld et al., 2024; Touvron et al., 2023) and Vision-Language Models (VLMs) Team et al. (2024a); Liu et al. (2023b); Deitke et al. (2024), giving rise to the paradigm of Vision-Language-Action models (VLAs) (Black et al., ; Brohan et al., 2022; Zheng et al., 2024; Zhao et al., 2025; Team et al., 2025; Li et al., 2024a; Qu et al., 2025; Kim et al., 2024). VLAs pretrain a VLM backbone on large-scale web data to capture rich semantic world knowledge, then fine-tune it for downstream robot control. While most share similar backbones, they differ in their action heads—employing flow matching(Wen et al., 2025b), diffusion(Wen et al., 2025a; Liu et al., 2024c), or advanced action tokenization schemes(Pertsch et al., 2025). While some adopt hierarchical designs, where a robotics-focused VLM (Bjorck et al., 2025; Li et al., 2025; Shentu et al., 2024) outputs intermediate representations for pre-trained, task-specific policies to improve generalization. However, a major bottleneck for these models is their heavy reliance on large amounts of robotics-specific data, often collected via teleoperation. In contrast, MolmoAct aims to leverage reasoning in space to train an action reasoning model that achieves competitive or superior performance with only a fraction of the data required by existing methods.
@@ 7 Conclusion
We introduced MolmoAct, a family of fully open action reasoning models that integrate perception, planning, and control by reasoning in space. By combining depth perception tokens, visual reasoning traces, and action prediction, MolmoAct produces explainable, spatially coherent behaviors. The behaviors can be executed directly, or steered via trajectory editing. Our evaluations across simulation and real-world settings demonstrate that MolmoAct consistently outperforms strong vision–language–action baselines, adapts efficiently to novel tasks and embodiments through lightweight fine-tuning, and generalizes robustly to out-of-distribution conditions. We release all model weights, code, and the MolmoAct Dataset to enable reproducibility and foster community-driven research toward building foundation models that transform perception into purposeful action through structured reasoning.
@@
  • Jason Lee: Led the Action Reasoning data curation for pre-training and post-training; Developed simulation evaluation infrastructure for pre-training and post-training; Led the real-world evaluations and simulation evaluations; contributed to real-world data collection.
  • Jiafei Duan: Led the project and ideated the core method design; proposed and curated the MolmoAct Dataset; spearheaded the paper’s writing; and designed and conducted both simulation and real-world evaluations, including ablation studies.
@@ Acknowledgment
  • We thank Christopher Clark, Abhay Deshpande, Yejin Kim, and Rohuan Tripathi for helpful research discussions and sharing of relevant findings across related projects.
  • We thank for David Ablbright, Crystal Nam, Kristin Cha, Sophie Lebrecht, Taira Anderson, Kyle Wiggers, Kelsey MacMillan, Katie Morigi, and Megan Bartot project management, support to robot room and publicity of MolmoAct

### v2

@@ 5.6 Steerability of MolmoAct
Evaluation Setups and Baselines. We aim to evaluate MolmoAct’s ability to steer robot actions, particularly when initial language instructions are ambiguous. Specifically, we investigate the effectiveness of different interaction mediums in guiding MolmoAct toward user-intended targets during task execution. For this purpose, we set up a pick_up_bowl task, post-training MolmoAct-7B-D and the baseline model ($\pi_{0}$-FAST) with 100 collected demonstrations, each annotated with two distinct language instructions: one specifying the clean bowl and the other the dirty bowl, as depicted in Figure 9. During evaluation, we first provide ambiguous instructions such as "pick up (the) bowl," prompting MolmoAct-7B-D to predict an initial trajectory towards one of the bowls. Subsequently, we test two steering methods: visual trace sketches to visually instruct the model toward the alternative bowl, and open-ended natural language instructions provided interactively by participants (N=10) that are different from the ground-truth instruction. For comparison, we also attempt to steer the actions of $\pi_{0}$-FAST by changing language instructions at test-time. Each model is evaluated in 15 trials, and the performance is evaluated according to the progression of the task. For more details about the setting, please refer to subsection D.7.
Evaluation Results. Based on our experiments, we observed that MolmoAct-7B-D is notably more steerable via visual trace inputs, achieving a success rate of 75%. Additionally, steering using visual traces significantly outperforms steering via open-ended natural language instructions by a margin of 33%. Lastly, we demonstrate that MolmoAct-7B-D exhibits superior instruction-following capabilities compared to the baseline model, $\pi_{0}$-FAST. Specifically, when steering robot actions using open-ended language instructions, MolmoAct surpasses $\pi_{0}$-FAST by a substantial margin of 29%, highlighting its enhanced instruction-following capabilities to user commands.
@@ 6.1 Generalist robot manipulation policies
Recent advances in robotic manipulation have shifted from training narrow, single-task specialists to learning from large, diverse datasets spanning many scenes, tasks, and embodiments(Berscheid et al., 2019; Brohan et al., 2022; Dasari et al., 2019; Ebert et al., 2021; Fang et al., 2023; Jang et al., 2022; Khazatsky et al., 2024; Mandlekar et al., 2018; Walke et al., 2023; Shafiullah et al., 2023). This shift has enabled policies that not only excel within their training distribution but also generalize to out-of-distribution scenes, environments, language instructions, and novel objects (Pumacay et al., 2024; Xie et al., 2024; Lin et al., 2024). Much of this progress has been fueled by Large Language Models (LLMs) (O’Neill et al., 2024; Achiam et al., 2023; Groeneveld et al., 2024; Touvron et al., 2023) and Vision-Language Models (VLMs) Team et al. (2024a); Liu et al. (2023b); Deitke et al. (2024), giving rise to the paradigm of Vision-Language-Action models (VLAs) (Black et al., ; Brohan et al., 2022; Zheng et al., 2024; Zhao et al., 2025; Team et al., 2025; Li et al., 2024a; Qu et al., 2025; Kim et al., 2024). VLAs pretrain a VLM backbone on large-scale web data to capture rich semantic world knowledge, then fine-tune it for downstream robot control. While most share similar backbones, they differ in their action heads—employing flow matching(Wen et al., 2025b), diffusion(Wen et al., 2025a; Liu et al., 2024c), or advanced action tokenization schemes(Pertsch et al., 2025). While some adopt hierarchical designs, where a robotics-focused VLM (Bjorck et al., 2025; Li et al., 2025; Shentu et al., 2024) outputs intermediate representations for pre-trained, task-specific policies to improve generalization. However, a major bottleneck for these models is their heavy reliance on large amounts of robotics-specific data, often collected via tele-operation. In contrast, MolmoAct aims to leverage reasoning in space to train an action reasoning model that achieves competitive or superior performance with only a fraction of the data required by existing methods.
@@ 7 Conclusion
We introduced MolmoAct, a family of fully open action reasoning models that integrate perception, planning, and control by reasoning in space. By combining depth perception tokens, visual reasoning traces, and action prediction, MolmoAct produces explainable, spatially coherent behaviors. The behaviors can be executed directly, or steered via trajectory editing. Our evaluations across simulation and real-world settings demonstrate that MolmoAct consistently outperforms strong vision–language–action baselines, adapts efficiently to novel tasks and embodiments through lightweight fine-tuning, and generalizes robustly to out-of-distribution conditions. We release all model weights, code, and data, including the MolmoAct Dataset, to enable reproducibility and foster community-driven research toward building foundation models that transform perception into purposeful action through structured reasoning.
@@
  • Jason Lee: Led the Action Reasoning data curation for pre-training, post-training and MolmoAct Dataset; Developed simulation evaluation infrastructure for pre-training and post-training; Led the real-world evaluations and simulation evaluations; contributed to real-world data collection and paper writing.
  • Jiafei Duan: Led the project and ideated the core method design; proposed and curated the MolmoAct Dataset; Led the paper’s writing; and designed and conducted both simulation and real-world evaluations, including ablation studies.
@@ Acknowledgment
  • We thank Christopher Clark, Abhay Deshpande, Yejin Kim, Mahtab Bigverdi, Joel Jang and Rohun Tripathi for helpful research discussions and sharing of relevant findings across related projects.
  • We thank for David Albright, Crystal Nam, Kristin Cha, Sophie Lebrecht, Taira Anderson, Kyle Wiggers, Kelsey MacMillan, Katie Morigi, and Megan Bartot for project management, support to robot room and publicity of MolmoAct